### v2

@@ Abstract
In response to the rapidly evolving nature of adversarial attacks against visual classifiers on a monthly basis, numerous defenses have been proposed to generalize against as many known attacks as possible. However, designing a defense method that generalizes to all types of attacks is not realistic because the environment in which defense systems operate is dynamic and comprises various unique attacks that emerge as time goes on. The defense system must gather online few-shot defense feedback to promptly enhance itself, leveraging efficient memory utilization. Therefore, we propose the first continual adversarial defense (CAD) framework that adapts to any attacks in a dynamic scenario, where various attacks emerge stage by stage. In practice, CAD is modeled under four principles: (1) continual adaptation to new attacks without catastrophic forgetting, (2) few-shot adaptation, (3) memory-efficient adaptation, and (4) high accuracy on both clean and adversarial images. We explore and integrate cutting-edge continual learning, few-shot learning, and ensemble learning techniques to qualify the principles. Experiments conducted on CIFAR-10 and ImageNet-100 validate the effectiveness of our approach against multiple stages of modern adversarial attacks and demonstrate significant improvements over numerous baseline methods. In particular, CAD is capable of quickly adapting with minimal feedback and a low cost of defense failure, while maintaining good performance against previous attacks. Our research sheds light on a brand-new paradigm for continual defense adaptation against dynamic and evolving attacks.
Keywords:
Adversarial Defense Continual Learning
@@
Figure 1: In a dynamic environment, the target model $f_{t}$ encounters a wide spectrum of attacks across various stages while being safeguarded by the defender. During the $i$-th stage, the defender leverages few-shot defense feedback, which is collected by the online defense system or provided by the security department and users of the target model, to adapt to emerging threats. Furthermore, the defender is required to efficiently manage memory utilization to store the cache, preserving insights gained from past attacks.
1 Introduction
In recent years, deep neural networks have been widely used in numerous vision tasks [29], leading to remarkable breakthroughs. However, many neural networks are vulnerable to adversarial attacks, which aim to deceive a classifier by adding subtle perturbations to input images, thus altering prediction results. This vulnerability seriously jeopardizes the reliability of deep neural networks, particularly in security- and trust-sensitive domains.
Researchers have proposed various general defense methods against adversarial attacks. As a standard defense method, adversarial training [44] aims to enhance the robustness of the target model by training it with adversarial examples. Another branch of adversarial defense involves purifying the data stream [41] to remove potential adversarial perturbations or noise that could deceive the model. However, models of adversarial training often exhibit reduced classification accuracy on adversarial examples and may sacrifice classification capacity on clean images. On the other hand, purification techniques may appear more reliable, but the denoising procedure can inadvertently smooth the high-frequency texture of clean images, potentially causing issues for the target model.
From an application perspective, it is not realistic to design a defense method that can generalize to all types of attacks, including unseen ones. The truth is that the increasing number of defense strategies naturally stimulates the development of new attack algorithms [32]. As a result, the defense system encounters a dynamic environment characterized by numerous unique attacks with higher success rates in both white-box and black-box settings [20]. The target model is threatened by many attackers with various attacks, just like the defense scenario against Deepfake [13].
Inspired by online learning, the crucial approach to addressing dynamic scenarios lies in embracing a perpetually evolving defense strategy. Consequently, deploying a defense system on the cloud that continuously gathers attack examples from newly emerged adversarial algorithms and promptly adapts to these attacks represents a promising method for responding to such scenarios. As depicted in Fig. 1, a variety of attacks are detected across different stages, with the defender adapting to emerging threats by utilizing few-shot defense feedback. In this paper, we employ adversarial examples with ground truths as the defense feedback, sourced either from the online defense system’s crawling efforts or provided by the security department and users of the target model in real-world scenarios. The defender undergoes training with an initial attack and shares the same training data and architecture as the target model. Meanwhile, attacks are initiated under the gray-box setting, wherein they possess knowledge of the classifier’s architecture and have access to training data, but remain unaware of the defense mechanism implemented.
@@ 1 Introduction
Taking real-world considerations into account, we propose four principles for the scenario: (1) Continual adaptation to new attacks without catastrophic forgetting. As defenders, it is crucial to adeptly adapt to a multitude of new attacks occurring at different stages while preserving knowledge gained from previous encounters. (2) Few-shot adaptation. The escalation in feedback signifies a greater potential for attacks on the target model. Therefore, we cannot tolerate an excessive number of attacks before adaptation. (3) Memory-efficient adaptation. Over time, a continuous influx of attacks leads to the accumulation of defense feedback, potentially posing memory constraints. In practical terms, we may not have sufficient memory capacity to withstand this pressure. (4) High accuracy in classifying both clean and adversarial images. A robust defense should not compromise the interests of those it protects. Hence, ensuring high classification accuracy on both clean and adversarial images is paramount.
@@
In this paper, we propose the Continual Adversarial Defense (CAD) framework, which aims to defend against evolving attacks in a stage-by-stage manner using few-shot defense feedback and efficient memory. In the first stage, we train an initial defense model using adversarial images from the initial attack. This model is specifically designed to classify images with adversarial noise, complementing the target model. Drawing inspiration from continual learning (CL), we expand the classification layer to adapt to the new attack as incremental classes. We then fine-tune this expanded layer using few-shot defense feedback. To address the over-fitting issue that arises from the few-shot fine-tuning process, we reserve embedding space in the defense model for future attacks. This is achieved by generating and assigning virtual prototypes, which help compress the embedding of previous attacks. To optimize memory usage in the data domain, we employ prototype augmentation which allows us to maintain the decision boundary of previous stages without the need to store any feedback explicitly. Simultaneously, we utilize a small model to ensemble the defense model with the target model by estimating reliable logits for input images, ensuring maximum classification accuracy for both clean and adversarial images.
@@
Based on our comprehensive experiments conducted on CIFAR-10 and ImageNet-100, the CAD framework demonstrates strong performance in defending against multi-stage attacks relying on few-shot feedback while maintaining high accuracy on clean images. Our main contributions can be summarized as follows:
@@
$\bullet$ To withstand the rapidly evolving adversarial attacks targeting image classifiers, we introduce, for the first time, a dynamic scenario featuring diverse attacks emerging at different stages to continuously enhance the defense. The defender must adeptly adapt to novel attacks using few-shot defense feedback and efficient memory management, while also retaining knowledge of past attacks and maintaining high performance on both clean and adversarial images.
@@
$\bullet$ We propose the Continual Adversarial Defense (CAD) framework that defends against attacks in the dynamic scenario under four practical principles: continual adaptation without catastrophic forgetting, few-shot adaptation, memory-efficient adaptation, and high accuracy in classifying both clean and adversarial images.We leverage cutting-edge techniques including continual learning, few-shot learning, non-exemplar class incremental learning, and ensemble learning to satisfy the principles.
@@
$\bullet$ Extensive experiments on CIFAR-10 and ImageNet-100 validate the effectiveness of CAD against multiple stages of 10 modern adversarial attacks using few-shot feedback and efficient memory and demonstrate significant improvements over 10 baseline methods.
2 Related Work
2.1 Adversarial Attacks
Adversarial attacks [9] attempt to fool a classifier into changing prediction results by attaching subtle perturbations to the input images while maintaining imperceptibility from human eyes. PGD [23] and BIM [18] generate adversarial examples through several iterations of perturbation and get significantly improved attack performance. DIM [40], TIM [7], and SIM [21] improve the transferability under the black-box setting and breach several defense techniques. Incorporating variance tuning, VIM [38] achieves a high attack rate against multiple advanced defense methods. In this paper, a wide spectrum of attacks is used to offense the target model at different stages under the gray-box setting.
2.2 Adversarial Defense
Adversarial training [23, 45] is the mainstream of adversarial defense that uses attacked samples to train a robust model. TRADES [44] characterize the trade-off between accuracy and robustness for classification problems via decomposing the robust error as the sum of the natural error and the boundary error. JEM [10] uses energy-based training of the joint distribution and improves the calibration and robustness of target models. Replacing the denoising diffusion probabilistic model with the elucidating diffusion model, DMAT [39] improves adversarial training and achieves new SOTA robust accuracy.
Another technical solution for protecting DNNs from adversarial attacks is adversarial purification [28, 12], which aims to purify the data stream before feeding it into the target model, functioning under the gray-box setting [32]. Based on an EBM trained with Denoising Score-Matching, Yoon et al. [41] propose Adaptive Denoising Purification (ADP) which conducts purification in a randomized purification scheme. Nie et al. [26] propose DiffPure that uses the forward and reverse processes of diffusion models to purify adversarial images.
Indeed, both adversarial training and purification methods have exhibited limitations in robustness, resulting in diminished classification performance on clean images. Moreover, the dynamic environment introduces a multitude of unique attacks, rendering a single defense inadequate. In this paper, we present the Continual Adversarial Defense (CAD) framework as a solution to effectively counter various attacks emerging stage by stage, while ensuring robust performance in classifying clean images.
2.3 Continual Learning
@@
Continual learning (CL) [22, 27] aims to learn from a sequence of new classes without forgetting old ones and attracts much attention to various computer vision tasks. Many works have been proposed for CL: by introducing different knowledge distillation (KD) losses to consolidate previous knowledge [27], by preserving a small number of old class data (exemplars) or augmented data (Zhu et al., 2021) to recall the old class knowledge [49], and by expanding the network capacity for new class data or freezing partial network parameters to keep the old class knowledge [47]. In recent years, some methods have aimed to solve the CL problem without depending on preserving data (called none-exemplar) [51], and some methods have attempted to learn new classes in a few-shot scenario [43] in which only a small number of new class data is gainable. In this paper, we convert the proposed defense scenario into a few-shot and non-exemplar CL setting, which necessitates the defense mechanism to utilize few-shot defense feedback and efficient memory for adaptation.
3 Threat Model and Defense Principles
@@
The defense system is deployed online and functions within a dynamic environment marked by a plethora of unique attacks, driven by the proliferation of defenses that naturally prompt the development of new attack algorithms. As time progresses, numerous new attacks, accompanied by attack examples, will emerge on the internet, aiming to exploit vulnerabilities in the defense system and achieve breakthroughs. In response to this reality, we simulate the dynamic environment and propose a practical threat model. In this model, distinct attacks are identified across various stages, requiring the defender to adapt to evolving threats through the utilization of defense feedback. Typically, defense feedback is obtained through crawling efforts of the online defense system or is provided by security services and users of the target model. This feedback is crucial as it informs and influences the defender’s response to emerging threats, often discovered during screening or cleanup efforts in the real world. In this paper, we use adversarial examples with ground truths as the defense feedback. We formulate this attack vs defense scenario as follows.
@@
Figure 2: Illustration of CAD. Regarding the classes with new attacks as incremental classes, we use continual adaptation to protect the target model while preventing catastrophic forgetting. In the 00-th stage, we reserve the embedding space for future attacks during training of the initial defense model $f_{0}$ to alleviate overfitting brought by future fine-tuning. Then, a weight estimator model $f_{p}$ is trained for model ensemble. In the $i$-th stage ($i>0$), we freeze the feature extractor $\varphi$ of defense model $f_{i}$ and fine-tune the expanded classifier $g_{c}^{i}$ using few-shot defense feedback. To efficiently use memory, we store class-wise prototypes in the prototype cache at each stage and use them in the fine-tuning process. After all, the target model $f_{t}$ and the scalable defense model $f_{i}$ are ensembled by the weight estimator model $f_{p}$ to maintain high accuracy in classifying clean and adversarial images.
@@
Our scenario is established on the $N$-way $K$-shot classification paradigm. Attacks are launched under the gray-box setting [32] in which the attacks know the architecture of the target model ${f}_{t}:\mathcal{X}\to\mathbb{R}^{N}$ which trained on $\mathcal{D}_{\mathrm{train}}$ and evaluated on $\mathcal{D}_{\mathrm{test}}$, but blind to the defense, where $\mathcal{X}$ is the image space. The defender is permitted to visit $\mathcal{D}_{\mathrm{train}}$ and has the full knowledge of the initial attack $A_{0}(\cdot)$ at the 0-th stage. At the $i$-th stage, where $i=1,2,\cdots,T$, a new attack $A_{i}(\cdot)$. is detected and the defender receives a set of feedback $\mathcal{A}_{\mathrm{train}}^{i}=\{(\mathbf{x}_{\mathrm{adv}}^{i},y)|\mathbf{x}_{\mathrm{adv}}^{i}=A_{i}(\mathbf{x}),(\mathbf{x},y)\in\mathcal{D}_{\mathrm{train}}\}$ containing $N\times K$ samples (i.e., $K$ samples for each of $N$ classes) and uses it to adapt to the new attack. Evaluations are conducted on $\mathcal{D}_{\mathrm{test}}$ and $\{\mathcal{A}_{\mathrm{test}}^{k}\}_{k=0,1,\dots,i}$ at each stage $i$, where $\mathcal{A}_{\mathrm{test}}^{i}=\{(\mathbf{x}_{\mathrm{adv}}^{i},y)|\mathbf{x}_{\mathrm{adv}}^{i}=A_{i}(\mathbf{x}),(\mathbf{x},y)\in\mathcal{D}_{\mathrm{test}}\}$.
@@
The increase in defense feedback directly correlates with the frequency of successful attacks on the target model. As defenders, we must proactively adapt the defense based on few-shot feedback to prevent potential security disasters that could arise from delayed action following an abundance of feedback.
@@
Over time, the ongoing influx of attacks results in accumulating defense feedback, potentially leading to memory constraints. In practical terms, the defender may not have sufficient memory space to accommodate this accumulation. Consequently, storing all received defense feedback for adaptation is impractical and should be avoided.
@@
High accuracy in classifying both clean and adversarial images.
The previous defense strategies have led to a sacrifice in the classification accuracy of clean images, resulting in performance degradation. This decline could potentially hurt crucial real-world business operations. As a result, a defender must prioritize maintaining performance on both clean and adversarial images.
4 Continual Adversarial Defense Framework
As shown in  Fig. 2, Continual Adversarial Defense (CAD) consists of continual adaptation to defend the target model while preventing catastrophic forgetting ( Sec. 4.1), embedding reservation to alleviate overfitting brought by the few-shot feedback ( Sec. 4.2), prototype augmentation for memory efficiency ( Sec. 4.3), and model ensemble to ensure the classification performance on both clean and adversarial images(Section 4.4).
4.1 Continual Adaptation to New Attacks
In response to Principle 1, we use continual adaptation to defend against new attacks while preventing catastrophic forgetting. In the 00-th stage, an initial defense model $f_{0}:\mathcal{X}\to\mathbb{R}^{N}$ which consists of a feature extractor $\varphi:\mathcal{X}\to\mathbb{R}^{d}$ and a classifier $g_{c}^{0}:\mathbb{R}^{d}\to\mathbb{R}^{N}$ is optimized under the full supervision using adversarial dataset $\mathcal{A}_{\mathrm{train}}^{0}=\{(\mathbf{x}_{\mathrm{adv}}^{0},y)|\mathbf{x}_{\mathrm{adv}}^{0}=A_{0}(\mathbf{x}),(\mathbf{x},y)\in\mathcal{D}_{\mathrm{train}}\}$. The defense model is designed to tackle adversarial examples, complementing the target model. In the $i$-th stage, the defense model ${f}_{i}:\mathcal{X}\to\mathbb{R}^{N}$ adapts to new attack using defense feedback $\mathcal{A}_{\mathrm{train}}^{i}$. Inspired by continual learning (CL) [27], we regard classes with the new attack as incremental classes and endow the defense model with scalability by classifier expansion.
Following [22], we expand the classifier from $g_{c}^{0}:\mathbb{R}^{d}\to\mathbb{R}^{N}$ to $g_{c}^{1}:\mathbb{R}^{d}\to\mathbb{R}^{N\times{2}}$ in the first stage, and expand the classifier $g_{c}^{i}$ from $\mathbb{R}^{d}\to\mathbb{R}^{N\times i}$ to $\mathbb{R}^{d}\to\mathbb{R}^{N\times{(i+1})}$ in the $i$-th stage. The parameter of the expanded classifier is composed of the parameter of the old classifier and the newly initialized parameter: $W^{i}=[W^{i-1},W^{new}]$, where $W^{i}=[\mathbf{w}^{i}_{1},\cdots,\mathbf{w}^{i}_{N_{i}}]$ is the parameter matrix of $g_{c}^{i}$ and $N_{i}=N\times(i+1)$ is number of classes in stage $i$ after expansion. Meanwhile, the ground truth of $\mathcal{A}_{\mathrm{train}}^{i}$ is rewritten corresponding to the incremental classes with the new attack  [^1] We omit the networks’ parameter in formulas:
@@ 4.1 Continual Adaptation to New Attacks
The initial defense model $f_{0}$ will be trained in Sec. 4.2. After this, we freeze the feature extractor $\varphi$ and fine-tune the part of the classifier expanded for the new attack $A_{i}$ using few-shot feedback $\mathcal{A}_{\mathrm{train}}^{i}$ in Sec. 4.3.
@@
4.2 Embedding Reservation for Few-Shot Adaptation
A problem brought by the few-shot feedback in Principle 2 is over-fitting which is common in few-shot learning [30]. To tackle this issue, we choose to reserve embedding space for future attacks by generating and assigning virtual prototypes, squeezing the embedding space of previous attacks.
First, before training the defense model $f_{0}$ at the 00-th stage, several virtual prototypes $P_{v}=[\mathbf{p}_{v}^{1},\cdots,\mathbf{p}_{v}^{V}]\in\mathbb{R}^{d\times V}$ are pre-assign in the classifier and treated as "virtual classes" [43], where $V=N\times T$ is the number of virtual classes, i.e. the reserved classes for future attacks. Therefore the output of current model is $f_{0}(\mathbf{x})=[W^{0},P_{v}]^{\top}\varphi(\mathbf{x})$.
Second, virtual instances are constructed by manifold mixup [35]:

### v3

@@ Abstract
In response to the rapidly evolving nature of adversarial attacks against visual classifiers on a monthly basis, numerous defenses have been proposed to generalize against as many known attacks as possible. However, designing a defense method that generalizes to all types of attacks is not realistic because the environment in which defense systems operate is dynamic and comprises various unique attacks that emerge as time goes on. A well-matched approach to the dynamic environment lies in a defense system that continuously collects adversarial data online to quickly improve itself. Therefore, we put forward a practical defense deployment against a challenging threat model and propose, for the first time, the Continual Adversarial Defense (CAD) framework that adapts to attack sequences under four principles: (1) continual adaptation to new attacks without catastrophic forgetting, (2) few-shot adaptation, (3) memory-efficient adaptation, and (4) high accuracy on both clean and adversarial data. We explore and integrate cutting-edge continual learning, few-shot learning, and ensemble learning techniques to qualify the principles. Extensive experiments validate the effectiveness of our approach against multiple stages of modern adversarial attacks and demonstrate significant improvements over numerous baseline methods. In particular, CAD is capable of quickly adapting with minimal budget and a low cost of defense failure while maintaining good performance against previous attacks. Our research sheds light on a brand-new paradigm for continual defense adaptation against dynamic and evolving attacks.
@@
Figure 1: In a dynamic environment, the target model $f_{t}$ faces a diverse range of attacks at different stages and is protected by the defender. During the $i$-th stage, the defense model leverages few-shot defense budget provided by the web crawler, security department, and users of the target model to adapt to new attacks. The defense model also needs to efficiently use memory to store the cache while retaining knowledge of past attacks.
Introduction
Adversarial attack (Madry et al. 2017) aims to deceive deep neural networks (DNNs) by adding subtle perturbations to input images, seriously jeopardizing the reliability of DNNs, particularly in domains sensitive to security and trust. In order to promote the robustness of DNN, adversarial training (Zhang et al. 2019) and purification techniques (Yoon, Hwang, and Lee 2021) are proposed to defend against adversarial attacks by one-shot training in which the model enters a static phase after a single defense training stage (Zhou and Hua 2024), but often exhibit reduced classification accuracy in adversarial examples and can sacrifice classification capacity in clean images.
Contrary to the one-shot defense assumption, the defense system deployed in real scenarios encounters a dynamic environment characterized by numerous unique attacks with higher aggressivity (Li et al. 2020). A well-matched approach to the dynamic environment lies in a constantly evolving defense strategy that continuously adapts to newly emerging adversarial algorithms. A concurrent continual defense work (Zhou and Hua 2024) treats attacks as a task sequence and tries to train a robust model through lifelong learning. However, poor performance in both adversarial and clean data hinders its deployment in reality and calls into question the reasonableness of the continual defense setting.
To eliminate the insufficiency of the above defense setting, we put forward a practical defense deployment and a challenging threat model. As depicted in Figure 1, the defense system is deployed in the cloud and continuously adapts to new emerging attacks from stage to stage using defense budget (i.e. adversarial examples and their ground-truths) sourced from the web crawler, security department, and users of the target model in real-world scenarios. In reality, a stage can consist of only a few samples from one or multiple attacks, and a new stage begins once the system gathers enough defense budget. When adapted to new emerging attacks, the defense system is permitted to use defense budget of previous stages cached in memory. The defense system also has access to an initial attack and shares the same training data and architecture as the target model. Meanwhile, attacks are initiated under the gray-box setting, wherein they possess knowledge of the classifier’s architecture and have access to training data, but remain unaware of the defense mechanism implemented.
@@ Introduction
Taking real-world considerations into account, we propose four principles for continual defense: (1) Continual adaptation to new attacks without catastrophic forgetting. As a defender, it is crucial to adeptly adapt to a multitude of new attacks occurring at different stages while preserving knowledge gained from previous encounters. (2) Few-shot adaptation. The escalation in the defense budget signifies a greater potential for attacks on the target model. Therefore, we cannot tolerate an excessive number of attacks before adaptation. (3) Memory-efficient adaptation. Over time, a continuous influx of attacks leads to the accumulation of defense budget, potentially posing memory constraints. In practical terms, we may not have sufficient memory capacity to withstand this pressure. (4) High accuracy in classifying clean and adversarial data. A robust defense should not compromise the interests of those it protects. Hence, ensuring high classification accuracy on both clean and adversarial data is paramount.
@@
In this paper, we propose the first Continual Adversarial Defense (CAD) framework, which aims to defend against evolving attacks in a stage-by-stage manner using few-shot defense budget and efficient memory. In the first stage, we use adversarial data from the initial attack to train an initial defense model which is specifically designed to classify adversarial data, complementing the target model. Drawing inspiration from continual learning (CL), we expand the classification layer to adapt to the new attack as incremental classes and then fine-tune this expanded layer using few-shot defense budget. To address the over-fitting issue that arises from the few-shot fine-tuning, we compress the embedding of previous attacks by generating and assigning virtual prototypes to reserve embedding space in the defense model for future attacks. To optimize memory usage in the data domain, we employ prototype augmentation which allows us to maintain the decision boundary of previous stages without the need to store any budget explicitly. After adaptation of one stage, we refine the classifier to prevent undesirable expansion. Simultaneously, a small model is utilized to ensemble the defense model with the target model by estimating reliable logits for input data, ensuring maximum classification accuracy for both clean and adversarial data.
@@
Based on comprehensive experiments conducted on CIFAR-10 and ImageNet-100, the CAD framework demonstrates strong performance in defending against multi-stage attacks relying on few-shot defense budget while maintaining high accuracy on clean data. To the best of our knowledge, this is the first time to discuss the concept and deployment scenario of continual adversarial defense.
Our main contributions can be summarized as follows:
@@
$\bullet$ To withstand the rapidly evolving adversarial attacks targeting DNNs, we put forward the deployment of continual defense strategy and propose a practical threat model in which diverse attacks emerge at different stages. The defense system must adapt well to these attacks using a few-shot defense budget and efficient memory management, while also retaining knowledge of past attacks and maintaining high performance on both clean and adversarial data.
@@
$\bullet$ We propose, for the first time, the Continual Adversarial Defense (CAD) framework that defends against attacks in the dynamic scenario under four practical principles: continual adaptation without catastrophic forgetting, few-shot adaptation, memory-efficient adaptation, and high accuracy in classifying both clean and adversarial data.We leverage cutting-edge techniques including continual learning, few-shot learning, non-exemplar class incremental learning, and ensemble learning to satisfy the principles.
@@
$\bullet$ Extensive experiments validate the effectiveness of CAD against multiple stages of modern adversarial attacks using few-shot budget and efficient memory and demonstrate significant improvements over baseline methods.
Related Work
Adversarial Defense
Researchers have proposed various general defense methods against adversarial attacks (Madry et al. 2017; Wu et al. 2021; Lin et al. 2019; Wang and He 2021). As a standard defense method, adversarial training (Zhang et al. 2020, 2019; Grathwohl et al. 2020; Wang et al. 2023b) aims to enhance the robustness of the target model by training it with adversarial examples. Another branch of adversarial defense involves purifying the data stream (Salman et al. 2020; Hill, Mitchell, and Zhu 2020; Yoon, Hwang, and Lee 2021; Nie et al. 2022) to remove potential adversarial perturbations or noise that could deceive the model. In fact, all of the above methods have shown limitations in robustness, resulting in a lower classification performance on both adversarial and clean data. In this paper, we present the first Continual Adversarial Defense (CAD) framework as a solution to effectively counter a wide spectrum of attacks emerging stage by stage, while ensuring robust performance in classifying clean data.
Continual Learning
@@
Continual learning (CL) (Liu et al. 2022; Kirkpatrick et al. 2017; Wang et al. 2022; Zhou et al. 2022) aims to learn from a sequence of new classes without forgetting old ones and attracts a lot of attention to various computer vision tasks. In recent years, some methods have been aimed at solving the CL problem without relying on preserving data (called no exemplar) (Zhu et al. 2022), and some methods have attempted to learn new classes in a few-shot scenario (Zhang et al. 2021) in which only a small number of new class data is gainable. Attempting to break the one-shot defense assumption, a concurrent work (Zhou and Hua 2024) is proposed to alleviate catastrophic forgetting in a simple continual defense scenario. In this paper, we convert the proposed defense scenario into a few-shot and non-exemplar CL setting, which necessitates the defense mechanism to utilize few-shot defense budget and efficient memory for adaptation.
Threat Model and Defense Principles
@@
The defense system is deployed online and protects the target model against a dynamic environment marked by a plethora of unique attacks. As time progresses, new attacks accompanied by poisoned examples will emerge on the Internet, aiming to exploit vulnerabilities in the target model and achieve breakthroughs. Due to the reality, we propose a practical threat model in which distinct attacks emerge across various stages, requiring the defense system to adapt to evolving threats through the utilization of the defense budget (i.e. adversarial examples and their ground-truths). Typically, a stage can consist of only a few samples from one or multiple attacks, and a new stage begins once the system gathers enough defense budget. The defense budget is crucial as it informs and influences the defender’s response to emerging threats, often discovered during screening or cleanup efforts in the real world, and is provided by web crawlers, security services, and users of the target model. The continual defense scenario is formulated as follows.
@@
Figure 2: Illustration of CAD. Regarding the classes with new attacks as incremental classes, we use continual adaptation to protect the target model while preventing catastrophic forgetting. In the 00-th stage, we reserve the embedding space for future attacks during training of the initial defense model $f_{0}$ to alleviate the overfitting brought by future fine-tuning. Then, a weight estimator model $f_{p}$ is trained for model ensemble. In the $i$-th stage ($i>0$), we freeze the feature extractor $\varphi$ of defense model $f_{i}$ and fine-tune the expanded classifier $g_{c}^{i}$ using few-shot defense budget. To efficiently use memory, we store class-wise prototypes in the prototype cache at each stage and use them in the fine-tuning process. After all, the classifier is refined, and the target model $f_{t}$ and the scalable defense model $f_{i}$ are ensembled by the weight estimator model $f_{p}$ to maintain high performance in classifying clean and adversarial data.
@@
Our scenario is established on the $N$-way $K$-shot classification paradigm. Attacks are launched under the gray-box setting (Taran et al. 2019) in which the attacks know the architecture of the target model ${f}_{t}:\mathcal{X}\to\mathbb{R}^{N}$ which trained on $\mathcal{D}_{\mathrm{train}}$ and evaluated on $\mathcal{D}_{\mathrm{test}}$, but blind to the defense, where $\mathcal{X}$ is the image space. The defender is permitted to visit $\mathcal{D}_{\mathrm{train}}$ and has the full knowledge of the initial attack $A_{0}(\cdot)$ at the 0-th stage. At the $i$-th stage, where $i=1,2,\cdots,T$, new attacks $A_{i}(\cdot)$. are detected and the defender receives a set of budget $\mathcal{A}_{\mathrm{train}}^{i}=\{(\mathbf{x}_{\mathrm{adv}}^{i},y)|\mathbf{x}_{\mathrm{adv}}^{i}=A_{i}(\mathbf{x}),(\mathbf{x},y)\in\mathcal{D}_{\mathrm{train}}\}$ containing $N\times K$ samples (i.e., $K$ samples for each of $N$ classes) and uses it to adapt to the new attack. Evaluations are conducted on $\mathcal{D}_{\mathrm{test}}$ and $\{\mathcal{A}_{\mathrm{test}}^{k}\}_{k=0,1,\dots,i}$ at each stage $i$, where $\mathcal{A}_{\mathrm{test}}^{i}=\{(\mathbf{x}_{\mathrm{adv}}^{i},y)|\mathbf{x}_{\mathrm{adv}}^{i}=A_{i}(\mathbf{x}),(\mathbf{x},y)\in\mathcal{D}_{\mathrm{test}}\}$.
@@
The increase in defense budget directly correlates with the frequency of successful attacks on the target model. The defense system must proactively adapt the defense based on few-shot budget to prevent potential security disasters that could arise from delayed action.
@@
Over time, the ongoing influx of attacks results in accumulating defense budgets, potentially leading to memory constraints. In practical terms, the defender may not have sufficient memory space to accommodate this accumulation. Consequently, storing all received defense budget for adaptation is impractical and should be avoided.
@@
High accuracy in classifying both clean and adversarial data.
The previous defense strategies have led to a sacrifice in the classification accuracy of clean data, resulting in performance degradation. This decline could potentially harm crucial real-world business operations. As a result, a defender must prioritize maintaining performance on both clean and adversarial data.
Continual Adversarial Defense Framework
As shown in Figure 2, Continual Adversarial Defense (CAD) framework consists of continual adaptation to defend the target model while preventing catastrophic forgetting, embedding reservation to alleviate overfitting brought by the few-shot budget, prototype augmentation for memory efficiency, and model ensemble to ensure the classification performance on both clean and adversarial data.
Continual Adaptation to New Attacks
In response to Principle 1, we use continual adaptation to defend against new attacks while preventing catastrophic forgetting. In the 00-th stage, an initial defense model $f_{0}:\mathcal{X}\to\mathbb{R}^{N}$ which consists of a feature extractor $\varphi:\mathcal{X}\to\mathbb{R}^{d}$ and a classifier $g_{c}^{0}:\mathbb{R}^{d}\to\mathbb{R}^{N}$ is optimized under full supervision using adversarial dataset $\mathcal{A}_{\mathrm{train}}^{0}=\{(\mathbf{x}_{\mathrm{adv}}^{0},y)|\mathbf{x}_{\mathrm{adv}}^{0}=A_{0}(\mathbf{x}),(\mathbf{x},y)\in\mathcal{D}_{\mathrm{train}}\}$. The defense model is designed to tackle adversarial examples, complementing the target model. In the $i$-th stage, the defense model ${f}_{i}:\mathcal{X}\to\mathbb{R}^{N}$ adapts to the new attack using the defense budget $\mathcal{A}_{\mathrm{train}}^{i}$.
Inspired by continual learning (CL) (Rebuffi et al. 2017), we regard the classes with the new attack as incremental classes and endow the defense model with scalability by classifier expansion. Due to the internal similarity of some adversarial attacks (Wang et al. 2023a), trained with one attack, the defense model can, to some extent, generalize to other attacks, resulting in some unnecessary class expansion. Old attacks occurring in new stages also lead to unnecessary adaptation. Therefore, we conduct model tests and filter out the useless data from the defense budget that can be correctly classified.
Following (Liu et al. 2022), we expand the classifier from $g_{c}^{0}:\mathbb{R}^{d}\to\mathbb{R}^{N}$ to $g_{c}^{1}:\mathbb{R}^{d}\to\mathbb{R}^{N\times{2}}$ in the first stage, and expand the classifier $g_{c}^{i}$ from $\mathbb{R}^{d}\to\mathbb{R}^{N\times i}$ to $\mathbb{R}^{d}\to\mathbb{R}^{N\times{(i+1})}$ in the $i$-th stage. The parameter of the expanded classifier is composed of the parameter of the old classifier and the newly initialized parameter: $W^{i}=[W^{i-1},W^{new}]$, where $W^{i}=[\mathbf{w}^{i}_{1},\cdots,\mathbf{w}^{i}_{N_{i}}]$ is the parameter matrix of $g_{c}^{i}$. Meanwhile, the ground truth of $\mathcal{A}_{\mathrm{train}}^{i}$ is rewritten corresponding to the incremental classes with the new attack  [^1] We omit the networks’ parameter in formulas:
@@
The initial defense model $f_{0}$ will be trained in the Embedding Reservation Section. After this, we freeze the feature extractor $\varphi$ and fine-tune the expanded classifier part for the new attack $A_{i}$ using the few-shot budget $\mathcal{A}_{\mathrm{train}}^{i}$ in the Prototype Augmentation Section.
@@
Embedding Reservation for Few-Shot Adaptation
A problem brought by the few-shot budget in Principle 2 is over-fitting which is common in few-shot learning (Snell, Swersky, and Zemel 2017). To tackle this issue, we choose to reserve embedding space for future attacks by generating and assigning virtual prototypes, squeezing the embedding space of previous attacks.
First, before training the defense model $f_{0}$ at the 00-th stage, several virtual prototypes $P_{v}=[\mathbf{p}_{v}^{1},\cdots,\mathbf{p}_{v}^{V}]\in\mathbb{R}^{d\times V}$ are pre-assign in the classifier and treated as ”virtual classes” (Zhang et al. 2021), where $V=N\times T$ is the number of virtual classes, i.e. the reserved classes for future attacks. Therefore the output of current model is $f_{0}(\mathbf{x})=[W^{0},P_{v}]^{\top}\varphi(\mathbf{x})$.
Second, virtual instances are constructed by manifold mixup (Verma et al. 2019):